# A Starting Point for Dynamic Community Detection with Leiden Algorithm

### v1

@@ A.4.3. Aggregation phase of our Parallel Leiden
Algorithm 7 Aggregation phase of our Parallel Leiden (Sahu, 2023).
1:$G^{\prime}$: Input/super-vertex graph
@@
In the algorithm, the process initiates by acquiring the offsets array for the community vertices CSR $G^{\prime}_{C^{\prime}}.offsets$ within lines 10-11. Initially, this involves tallying the number of vertices associated with each community using countCommunityVertices(), followed by executing an exclusive scan operation on the array. Subsequently, within lines 12-13, we concurrently traverse all vertices, atomically assigning vertices linked with each community into the community graph CSR $G^{\prime}_{C^{\prime}}$. Moving forward, we determine the offsets array for the super-vertex graph CSR by approximating the degree of each super-vertex within lines 15-16. This involves computing the cumulative degree of each community utilizing communityTotalDegree() followed by executing an exclusive scan on the array. Consequently, the super-vertex graph CSR displays intervals between the edges and weights array of each super-vertex. Subsequent to this, within lines 18-24, we iterate over all communities $c\in[0,|\Gamma|)$. Here, we include all communities $d$ (alongside their respective edge weight $w$) connected to each vertex $i$ belonging to community $c$ (through scanCommunities() as defined in Algorithm 5) into the per-thread hashtable $H_{t}$. Upon $H_{t}$ containing all communities (with their weights) linked to community $c$, we atomically append them as edges to super-vertex $c$ within the super-vertex graph $G^{\prime\prime}$. Ultimately, in line 25, we return $G^{\prime\prime}$.
A.5. Updating vertex/community weights
@@
Algorithm 8 Updating vertex/community weights in parallel (Sahu, 2024).
1:$G^{t}$: Current input graph

### v4

@@ A.3.3. Aggregation phase of our Parallel Leiden
In the algorithm, the process begins by obtaining the offsets array for the community vertices in the CSR format, denoted as $G^{\prime}_{C^{\prime}}.offsets$, within lines 10 to 11. This starts with counting the number of vertices in each community using countCommunityVertices(), followed by performing an exclusive scan on the resulting array. Next, within lines 12 to 13, we concurrently traverse all vertices, atomically placing the vertices associated with each community into the community graph CSR $G^{\prime}_{C^{\prime}}$. Following this, the offsets array for the super-vertex graph CSR is computed by estimating the degree of each super-vertex within lines 15 to 16. This involves calculating the total degree of each community using communityTotalDegree(), followed by another exclusive scan. As a result, the super-vertex graph CSR is organized with intervals for the edges and weights array of each super-vertex. Then, in lines 18 to 24, we iterate over all communities $c\in[0,|\Gamma|)$ in parallel using dynamic loop scheduling, with a chunk size of $2048$ for both Static Leiden and a chunk size of $32$ for ND, DS, and DF Leiden. During this phase, all communities $d$ (and their respective edge weights $w$) connected to each vertex $i$ in community $c$ are included (via scanCommunities(), described in Algorithm 6) in the per-thread hashtable $H_{t}$. Once $H_{t}$ contains all the connected communities and their weights, they are atomically added as edges to super-vertex $c$ in the super-vertex graph $G^{\prime\prime}$. Finally, in line 25, we return the super-vertex graph $G^{\prime\prime}$.
Algorithm 8 Aggregation phase of our Parallel Leiden (Sahu et al., 2024a).
1:$G^{\prime}(V^{\prime},E^{\prime})$: Input/super-vertex graph
@@
A.4. Updating vertex/community weights
@@
Algorithm 9 Updating vertex/community weights in parallel.
1:$G^{t}(V^{t},E^{t})$: Current input graph
@@
A.5. Performance Comparison on Real-world dynamic graphs
We also evaluate the performance of our parallel implementations of Static, ND, DS, and DF Leiden on real-world dynamic graphs listed in Table 3. These evaluations are performed on batch updates ranging from $10^{-5}|E_{T}|$ to $10^{-3}|E_{T}|$ in multiples of $10$. For each batch size, as described in Section 5.1.4, we load $90\%$ of the graph, add reverse edges to ensure all edges are undirected, and then load $B$ edges (where $B$ is the batch size) consecutively in $100$ batch updates. Figure 11(a) shows the overall runtime of each approach across all graphs for each batch size, while Figure 11(b) depicts the overall modularity of the obtained communities. Additionally, Figures 11(c) and 11(d) present the mean runtime and modularity of the communities obtained with each approach on individual dynamic graphs in the dataset.
Figure 11(a) illustrates that ND Leiden is, on average, 1.09×1.09$\times$1.09 × faster than Static Leiden for batch updates ranging from $10^{-5}|E_{T}|$ to $10^{-3}|E_{T}|$. In comparison, DS and DF Leiden exhibit average speedups of 1.20×1.20$\times$1.20 × and 1.33×1.$3$, respectively, over Static Leiden for the same batch updates. We now explain why ND, DS, and DF Leiden achieve only minor speedups over Static Leiden. Our experiments indicate that only about $20\%$ of the overall runtime of Static Leiden is spent in the local-moving phase of the first pass. Consequently, the speedup achieved by ND, DS, and DF Leiden relative to Static Leiden is limited. Furthermore, our observations suggest that while DF Leiden can reduce the time spent in the local-moving phase slightly more than ND Leiden, it incurs increased runtimes in the refinement and aggregation phases of the algorithm. This is likely because ND Leiden is able to optimize clusters more effectively than DF Leiden, which accelerates the refinement and aggregation phases for ND Leiden, albeit by a small margin. In addition, we observe that only about $20\%$ of the overall runtime of Static Leiden is spent in the local-moving phase of the first pass. This explains the lower speedup of DF Leiden compared to ND Leiden on larger batch updates of real-world dynamic graphs.
(a) Overall Runtime
(b) Overall Modularity of communities obtained
(c) Runtime on each dynamic graph
(d) Modularity in communities obtained on each dynamic graph